OpenStudy (anonymous):

solve by elimination method. 0.3x-0.2y=4 0.5x+0.3y=-7/17

OpenStudy (amistre64):

....... fractions and decies eh

OpenStudy (amistre64):

0.3x-0.2y=4 0.5x+0.3y=-7/17 ; *10 the whole shebang 3x- 2y=40 5x+ 3y=-70/17 <- *17 3x- 2y = 40 85x+51y=-70

OpenStudy (amistre64):

3x- 2y = 40 <- * -85/3 i spose 85x+51y=-70

OpenStudy (amistre64):

-85x +(85)2/3y = -(85)40/3 85x +51(3)/3y= -70(3)/3 --------------------------- 323/3y = -3610/3 y = -3610/323 thats messy

OpenStudy (anonymous):

i got that far, but am lost right after that

OpenStudy (dumbcow):

Instead of substituting back in y to get x, you could try multiply top by 51 and bottom by 2 153x -102y = 2040 170x +102y = -140 __________________ 323x = 1900 x = 1900/323